OpenStudy (anonymous):

In a scale drawing, a 343 foot antenna tower is drawn 7 inches high. The building next to it is drawn 5 inches high. How tall is the actual building?

OpenStudy (amistre64):

h varies directly to scale 343 = k7 k = 343/7 = 49 h = 49s h = 49(5) = 245

OpenStudy (anonymous):

got it thank you why do people wnat medals so bad?? Im knew to this